OpenStudy (anonymous):

the volumes of two similar pyramids are 27 and 64. If the smaller has lateral area of 18, find the lateral area of the larger

OpenStudy (anonymous):

k = ratio of the sides of the bases = ratio of the heights. k^2 = ratio of the areas k^3 = ration of the volumes

OpenStudy (anonymous):

volume of a pyramid = (1/3)(base)(height) = (1/3)bh

OpenStudy (anonymous):

ratio of the volumes is 27:64, so k = 3/4 k^2 = 9/16 so if the smaller pyramid has area of 18, then the larger pyramid has area of 32

OpenStudy (anonymous):

get it?